# Intersections of blocks of cyclotomic Hecke algebras

### v1

@@ 3. Blocks in the exceptional types
It is our understanding that in [31, §8] it is shown for types $G_{2}$ and $F_{4}$ that the block sizes of the specialised cyclotomic algebras do match up as predicted by Conjecture 2.3, but without making explicit a choice of HC-parametrisation and (hence) also without checking that the images of the blocks under $\Psi_{{\mathbf{L}},\lambda},\Psi_{{\mathbf{M}},\mu}$ do actually agree. We check this latter property using the data and the functionality of the Chevie-system [28].
@@ 3.1. Type $G_{2}$
We start with type $G_{2}$. The block sizes for the relevant algebras have already been determined in [31, 8.6]. Here we also consider the intersection of the corresponding sets of unipotent characters. There are two pairs of characters of $G_{2}(q)$ not distinguished by their degrees, namely the principal series characters $\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime}$ and the cuspidal characters $G_{2}[\theta],G_{2}[\theta^{2}]$ of $G_{2}(q)$, where as customary $\theta$ denotes a primitive third root of unity. For these, there exist several possible Harish-Chandra parametrisations. Now it turns out that $\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime}$ in all cases either lie in the same block or they both occur in a singleton, so for the verification of Conjecture 2.3 no ambiguity arises. On the other hand, for $G_{2}[\theta],G_{2}[\theta^{2}]$ a parametrisation via endomorpism algebras is fixed by [19, (7.3)] for $d=6$. Hence the only relevant choice is for $d=3$. We present a compatible HC-parametrisation in Table 1. Here the characters of ${\mathcal{H}}(W(G_{2}),x)$ are identified through their standard Chevie-names, the (linear) character for cyclotomic Hecke algebras for cyclic groups by the corresponding parameter as in ()($\dagger$)( † ). With this, we have:
@@
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6}\}\\ (1,3)&\{\phi_{1,0},\phi_{1,6},\phi_{2,2}\}\\ (1,6)&\{\phi_{1,0},\phi_{1,6},\phi_{2,1}\}\\ (3,6)&\{\phi_{1,0},G_{2}[\theta]\},\{\phi_{1,6},G_{2}[\theta^{2}]\}\\ \end{array}$
@@
The relative Weyl groups are cyclic or of type $G_{2}$. The parameters for the cyclotomic Hecke algebra are known by the work of Lusztig or predicted in [5, Bem. 2.7] (see also the table compiled in [31, 8.6]). The decomposition matrices for specialisations of ${\mathcal{H}}(W(G_{2}),x)$ are given in [15, Tab. 7.3], the ones for the cyclic case were described in Section 2.3. The claim follows by direct computation. ∎
@@ 3.2. Type ${}^{3}\!D_{4}$
All unipotent characters of the triality groups ${}^{3}\!D_{4}(q)$ are uniquely determined by their degree polynomial and hence ()(*)( ∗ ) determines a unique complete HC-parametrisation.
@@
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6}\}\\ (1,3)&\{\phi_{1,0},\phi_{1,6},\phi_{2,2}\},\{\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{2,1}\}\\ (1,6)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6},\phi_{2,2}\}\\ (1,12)&\{\phi_{1,0},\phi_{1,6},\phi_{2,1}\}\\ (3,6)&\{\phi_{1,0},\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime},\phi_{1,6},\phi_{2,2},{}^{3}\!D_{4}[1]\}\\ (3,12)&\{\phi_{1,0},\phi_{1,6},\phi_{2,1}\}\\ \end{array}$
@@
The relative Weyl groups are described in [6, Tab. 3], the parameters for the corresponding cyclotomic Hecke algebras are predicted in [5, Folg. 5.6] (see also the table compiled in [31, 8.8.1]). The blocks for non-cyclic $W$ are again to be found in [15, Tab. 7.4] for type $W(G_{2})$, and have been worked out in [12, 3.3.1] for the primitive complex reflection group $G_{4}$, respectively. ∎
@@ 3.3. Type $F_{4}$
We now turn to type $F_{4}$. Again, the block sizes for the relevant algebras have already been determined by Trinh and Xue [31, 8.7]. Here, we also compare the actual intersections of Harish-Chandra series. This group is particularly interesting as there are ten pairs of unipotent characters of $F_{4}(q)$ having equal degree, so for all of those are there potential choices in the HC-parametrisation.
@@
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{\phi_{1,0},\phi_{1,12}^{*},\phi_{1,24},\phi_{2,4}^{*},\phi_{2,16}^{*},\phi_{9,2},\phi_{9,6}^{*},\phi_{9,10},\phi_{6,6}^{*},\phi_{8,3}^{*},\phi_{8,9}^{*}\},\\ &\{\phi_{4,1},\phi_{4,7}^{*},\phi_{4,13}\},\{B_{2}[.1^{2}],B_{2}[.2],B_{2}[1^{2}.],B_{2}[2.]\}\\ (1,3)&\{\phi_{1,0},\phi_{1,12}^{*},\phi_{1,24},\phi_{2,4}^{*},\phi_{2,16}^{*},\phi_{4,8}\},\{\phi_{4,1},\phi_{4,7}^{*},\phi_{4,13},\phi_{8,3}^{*},\phi_{8,9}^{*},\phi_{16,5}\}\\ (1,4)&\{\phi_{1,0},\phi_{1,24},\phi_{4,8},\phi_{9,2},\phi_{9,10},\phi_{6,6}^{\prime},\phi_{12,4},\phi_{4,1},\phi_{4,13}\},\\ &\{\phi_{2,4}^{\prime},\phi_{2,16}^{\prime},\phi_{4,7}^{\prime}\},\{\phi_{2,4}^{\prime\prime},\phi_{2,16}^{\prime\prime},\phi_{4,7}^{\prime\prime}\},\{B_{2}[.1^{2}],B_{2}[1.1],B_{2}[2.]\}\\ (1,6)&\{\phi_{1,0},\phi_{1,24},\phi_{2,4}^{*},\phi_{2,16}^{*},\phi_{9,6}^{*},\phi_{12,4},\phi_{8,3}^{*},\phi_{8,9}^{*}\},\\ &\{B_{2}[.1^{2}],B_{2}[.2],B_{2}[1^{2}.],B_{2}[2.]\}\\ (1,12)&\{\phi_{1,0},\phi_{1,24},\phi_{6,6}^{\prime\prime},\phi_{4,1},\phi_{4,13}\},\{B_{2}[.1^{2}],B_{2}[1.1],B_{2}[2.]\}\\ (3,4)&\{\phi_{1,0},\phi_{1,24},\phi_{4,8},\phi_{4,1},\phi_{4,13},F_{4}^{2}[1]\},\{\phi_{2,4}^{\prime},\phi_{2,16}^{\prime},\phi_{4,7}^{\prime}\},\{\phi_{2,4}^{\prime\prime},\phi_{2,16}^{\prime\prime}\phi_{4,7}^{\prime\prime}\},\\ (3,6)&\{\phi_{1,0},\phi_{2,16}^{*},\phi_{8,3}^{*},F_{4}[\theta^{2}]\},\{\phi_{1,24},\phi_{2,4}^{*},\phi_{8,9}^{*},F_{4}[\theta]\},\\ (3,12)&\{\phi_{1,0},\phi_{4,13},F_{4}[\theta]\},\{\phi_{1,24},\phi_{4,1},F_{4}[\theta^{2}]\}\\ (4,8)&\{\phi_{1,0},\phi_{9,10},F_{4}[i]\},\{\phi_{1,24},\phi_{9,2},F_{4}[-i]\}\\ (4,12)&\{\phi_{1,0},\phi_{1,24},B_{2}[1.1]\},\{\phi_{4,1},B_{2}[.1^{2}],F_{4}[-i]\},\{\phi_{4,13},B_{2}[2.],F_{4}[i]\}\\ (8,12)&\{\phi_{1,0},F_{4}[-i]\},\{\phi_{1,24},F_{4}[i]\}\\ \end{array}$
@@ 3.4. Type $E_{6}$
For $G$ of type $E_{6}$, again most block sizes for the relevant algebras have been determined in [31, 8.7]. Here, we also compare the blocks to the intersections of Harish-Chandra series. The only unipotent characters not distinguished by their degree are the two cuspidal ones, so only for those is there a choice in the HC-parametrisation.
@@ 4.1. A generalised conjecture
and an action of $N_{\operatorname{GL}(V)}(W\varphi)$ on ${\operatorname{Uch}}({\mathbb{G}})$ (see [20] for the non-rational Coxeter groups, [22] and [24, Thm 8.4] for the non-real cases, and also [8] for explicit tables for primitive $W$). If $W$ is a Weyl group, ${\operatorname{Uch}}({\mathbb{G}})$ are just the (labels of the) unipotent characters of an associated finite reductive group of rational type $W\varphi$ together with their degree polynomials. For any root of unity $\zeta$ and $\Phi:=x-\zeta$, there is defined a subset of $\Phi$-cuspidal elements of ${\operatorname{Uch}}({\mathbb{G}})$. A (unipotent) $\Phi$-cuspidal pair of ${\mathbb{G}}$ is a pair $({\mathbb{L}},\lambda)$ where ${\mathbb{L}}$ is a $\Phi$-split Levi of ${\mathbb{G}}$ and $\lambda\in{\operatorname{Uch}}({\mathbb{L}})$ is $\Phi$-cuspidal. For $({\mathbb{L}},\lambda)$ a $\Phi$-cuspidal pair let $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ be the stabiliser of $\lambda$ in $W_{\mathbb{G}}({\mathbb{L}})$ (see [8, Ax. 4.3]). Attached to $({\mathbb{L}},\lambda)$ is also a Hecke algebra ${\mathcal{H}}(W_{\mathbb{G}}({\mathbb{L}},\lambda),x)$ in one parameter $x$ (a specialisation of the generic Hecke algebra of $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ for parameters specified in [22, 21, 8]) naturally isomorphic (via the specialisation $x\mapsto\zeta$ and Tits’ deformation theorem) to the group algebra of $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ over a splitting field. This Hecke algebra should carry a canonical symmetrising form with associated Schur elements denoted $S_{\chi}$, see Section 2.1. According to [8, Ax. 4.31], there is a partition
@@
such that the exact analogue of ()(*)( ∗ ) holds. It thus makes sense to propose the following, where $({\mathbb{M}},\mu)$ is a $\Phi^{\prime}=x-\zeta^{\prime}$-cuspidal pair of ${\mathbb{G}}$:
@@
It seems reasonable to expect that arguments similar to those used in [31, §6] might be applicable to show this for the infinite series of imprimitive spetsial groups.
@@ 4.2. The non-crystallographic Coxeter groups
When $m$ is even, the case when $d=1$ and $e\geq 3$ is entirely similar to the above, so it remains to deal with the case $d=1,e=2$. Here, by [15, Tab. 7.2] the Iwahori–Hecke algebra ${\mathcal{H}}(I_{2}(m),(x,-1))$ specialised at $x=-1$ has just one non-singleton block, containing the four linear characters. By Ennola duality again, the Hecke algebra ${\mathcal{H}}(I_{2}(m),(x,1))$ specialised at $x=1$ also has just one non-singleton block, again consisting of the of the four linear characters. By the description in [22, Def. 6.3] the linear characters in both cases label the same four unipotent characters, confirming Conjecture 4.1 in this case. In fact, this case also follows from the analogue of Lemma 2.7(b). ∎
@@
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,3)&\{\phi_{1,0},\phi_{4,3},\phi_{5,2}\},\{\phi_{1,15},\phi_{4,4},\phi_{5,5}\}\\ (1,5^{\prime})&\{\phi_{1,0},\phi_{3,6},\phi_{4,3}\},\{\phi_{1,15},\phi_{3,1},\phi_{4,4}\}\\ (1,10^{\prime})&\{\phi_{1,0},\phi_{1,15},\phi_{3,1},\phi_{3,6}\},\\ &\{H_{2}[\zeta_{5}^{3}]{:}1^{2},H_{2}[\zeta_{5}^{3}]{:}2\},\{H_{2}[\zeta_{5}^{2}]{:}1^{2},H_{2}[\zeta_{5}^{2}]{:}2\}\\ (3,5^{\prime})&\{\phi_{1,0},\phi_{4,3}\},\{\phi_{1,15},\phi_{4,4}\}\\ (3,6)&\{\phi_{1,0},\phi_{5,5}\},\{\phi_{1,15},\phi_{5,2}\}\\ (5^{\prime},5^{\prime\prime})&\{\phi_{1,0},H_{2}[\zeta_{5}^{3}]{:}1^{2}\},\{\phi_{1,15},H_{2}[\zeta_{5}^{2}]{:}2\},\{\phi_{4,3},H_{2}[\zeta_{5}^{2}]{:}1^{2}\},\{\phi_{4,4},H_{2}[\zeta_{5}^{3}]{:}2\}\\ (5^{\prime},10^{\prime})&\{\phi_{1,0},H_{2}[\zeta_{5}^{2}]{:}2\},\{\phi_{1,15},H_{2}[\zeta_{5}^{3}]{:}1^{2}\},\{\phi_{3,1},H_{2}[\zeta_{5}^{2}]{:}1^{2}\},\{\phi_{3,6},H_{2}[\zeta_{5}^{3}]{:}2\}\\ (5^{\prime},10^{\prime\prime})&\{\phi_{1,0},H_{2}[\zeta_{5}^{2}]{:}1^{2}\},\{\phi_{1,15},H_{2}[\zeta_{5}^{3}]{:}2\},\{H_{2}[\zeta_{5}^{3}]{:}1^{2},H_{2}[\zeta_{5}^{2}]{:}2\}\\ \end{array}$
@@ 4.3. Verifying the spetsial conjecture
$\begin{array}[]{c|cc}(d,e)&\text{intersections}\\ \hline\cr(1,2)&\{[\phi_{1,0},\phi_{1,4},\phi_{1,8},\phi_{2,3},\phi_{2,1},\phi_{3,2}\},\{Z_{3}[2],Z_{3}[1^{2}]\}\\ (1,3^{\prime})&\{\phi_{1,0},\phi_{1,8}\},\{\phi_{2,5},\phi_{2,1}\}\\ (1,4)&\{\phi_{1,0},\phi_{2,5},\phi_{3,2}\}\\ (1,6^{\prime})&\{\phi_{1,0},\phi_{1,4},\phi_{2,1}\},\{Z_{3}[2],Z_{3}[1^{2}]\}\\ (3^{\prime},3^{\prime\prime})&\{\phi_{1,0},Z_{3}[1^{2}]\},\{\phi_{2,5},Z_{3}[2]\}\\ (3^{\prime},4)&\{\phi_{1,0},\phi_{2,5}\}\\ (3^{\prime},6^{\prime})&\{\phi_{1,0},Z_{3}[2]\},\{\phi_{2,1},Z_{3}[1^{2}]\}\\ (3^{\prime},6^{\prime\prime})&\{\phi_{1,0},Z_{3}[1^{2}]\},\{\phi_{1,8},Z_{3}[2]\}\\ \end{array}$

### v2

@@ 3. Blocks in the exceptional types
It is our understanding that in [32, §8] it is shown for types $G_{2}$ and $F_{4}$ that the block sizes of the specialised cyclotomic algebras do match up as predicted by Conjecture 2.3, but without making explicit a choice of HC-parametrisation and (hence) also without checking that the images of the blocks under $\Psi_{{\mathbf{L}},\lambda},\Psi_{{\mathbf{M}},\mu}$ do actually agree. We check this latter property using the data and the functionality of the Chevie-system [28].
@@ 3.1. Type $G_{2}$
We start with type $G_{2}$. The block sizes for the relevant algebras have already been determined in [32, 8.6]. Here we also consider the intersection of the corresponding sets of unipotent characters. There are two pairs of characters of $G_{2}(q)$ not distinguished by their degrees, namely the principal series characters $\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime}$ and the cuspidal characters $G_{2}[\theta],G_{2}[\theta^{2}]$ of $G_{2}(q)$, where as customary $\theta$ denotes a primitive third root of unity. For these, there exist several possible Harish-Chandra parametrisations. Now it turns out that $\phi_{1,3}^{\prime},\phi_{1,3}^{\prime\prime}$ in all cases either lie in the same block or they both occur in a singleton, so for the verification of Conjecture 2.3 no ambiguity arises. On the other hand, for $G_{2}[\theta],G_{2}[\theta^{2}]$ a parametrisation via endomorpism algebras is fixed by [19, (7.3)] for $d=6$. Hence the only relevant choice is for $d=3$. We present a compatible HC-parametrisation in Table 1. Here the characters of ${\mathcal{H}}(W(G_{2}),x)$ are identified through their standard Chevie-names, the (linear) character for cyclotomic Hecke algebras for cyclic groups by the corresponding parameter as in $(\dagger)$. With this, we have:
@@
[…]
@@
The relative Weyl groups are cyclic or of type $G_{2}$. The parameters for the cyclotomic Hecke algebra are known by the work of Lusztig or predicted in [5, Bem. 2.7] (see also the table compiled in [32, 8.6]). The decomposition matrices for specialisations of ${\mathcal{H}}(W(G_{2}),x)$ are given in [15, Tab. 7.3], the ones for the cyclic case were described in Section 2.3. The claim follows by direct computation. ∎
@@ 3.2. Type $^{3}\!D_{4}$
All unipotent characters of the triality groups $^{3}\!D_{4}(q)$ are uniquely determined by their degree polynomial and hence ()(*) determines a unique complete HC-parametrisation.
@@
[…]
@@
The relative Weyl groups are described in [6, Tab. 3], the parameters for the corresponding cyclotomic Hecke algebras are predicted in [5, Folg. 5.6] (see also the table compiled in [32, 8.8.1]). The blocks for non-cyclic $W$ are again to be found in [15, Tab. 7.4] for type $W(G_{2})$, and have been worked out in [12, 3.3.1] for the primitive complex reflection group $G_{4}$, respectively. ∎
@@ 3.3. Type $F_{4}$
We now turn to type $F_{4}$. Again, the block sizes for the relevant algebras have already been determined by Trinh and Xue [32, 8.7]. Here, we also compare the actual intersections of Harish-Chandra series. This group is particularly interesting as there are ten pairs of unipotent characters of $F_{4}(q)$ having equal degree, so for all of those are there potential choices in the HC-parametrisation.
@@
[…]
@@ 3.4. Type $E_{6}$
For $G$ of type $E_{6}$, again most block sizes for the relevant algebras have been determined in [32, 8.7]. Here, we also compare the blocks to the intersections of Harish-Chandra series. The only unipotent characters not distinguished by their degree are the two cuspidal ones, so only for those is there a choice in the HC-parametrisation.
@@ 4.1. A generalised conjecture
and an action of $N_{\operatorname{GL}(V)}(W\varphi)$ on ${\operatorname{Uch}}({\mathbb{G}})$ (see [20] for the non-rational Coxeter groups, [22] and [24, Thm 8.4] for the non-real cases, and also [8] for explicit tables for primitive $W$). If $W$ is a Weyl group, ${\operatorname{Uch}}({\mathbb{G}})$ are just the (labels of the) unipotent characters of an associated finite reductive group of rational type $W\varphi$ together with their degree polynomials. For any root of unity $\zeta$ and $\Phi:=x-\zeta$, there is defined a subset of $\Phi$-cuspidal elements of ${\operatorname{Uch}}({\mathbb{G}})$. A (unipotent) $\Phi$-cuspidal pair of $\mathbb{G}$} is a pair $({\mathbb{L}},\lambda)$ where $\mathbb{L}$} is a $\Phi$-split Levi of $\mathbb{G}$} and $\lambda\in{\operatorname{Uch}}({\mathbb{L}})$ is $\Phi$-cuspidal. For $({\mathbb{L}},\lambda)$ a $\Phi$-cuspidal pair let $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ be the stabiliser of $\lambda$ in $W_{\mathbb{G}}({\mathbb{L}})$ (see [8, Rem. 4.4]). It turns out by inspection [8, Rem. 4.15] that this is again a complex reflection group. Attached to $({\mathbb{L}},\lambda)$ is also a Hecke algebra ${\mathcal{H}}(W_{\mathbb{G}}({\mathbb{L}},\lambda),x)$ in one parameter $x$ (a specialisation of the generic Hecke algebra of $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ for parameters specified in [22, 21, 8]) naturally isomorphic (via the specialisation $x\mapsto\zeta$ and Tits’ deformation theorem) to the group algebra of $W_{\mathbb{G}}({\mathbb{L}},\lambda)$ over a splitting field. This Hecke algebra should carry a canonical symmetrising form with associated Schur elements denoted $S_{\chi}$, see Section 2.1. According to [8, Ax. 4.31], there is a partition
@@
such that the exact analogue of ()(*) holds. It thus makes sense to propose the following, where $({\mathbb{M}},\mu)$ is a $\Phi^{\prime}=x-\zeta^{\prime}$-cuspidal pair of $\mathbb{G}$}:
@@
It seems reasonable to expect that arguments similar to those used in [32, §6] might be applicable to show this for the infinite series of imprimitive spetsial groups.
@@ 4.2. The non-crystallographic Coxeter groups
When $m$ is even, the case when $d=1$ and $e\geq 3$ is entirely similar to the above, so it remains to deal with the case $d=1,e=2$. Here, by [15, Tab. 7.2] the Iwahori–Hecke algebra ${\mathcal{H}}(I_{2}(m),(x,-1))$ specialised at x=1x=-1 has just one non-singleton block, containing the four linear characters. By Ennola duality again, the Hecke algebra ${\mathcal{H}}(I_{2}(m),(x,1))$ specialised at $x=1$ also has just one non-singleton block, again consisting of the of the four linear characters. By the description in [22, Def. 6.3] the linear characters in both cases label the same four unipotent characters, confirming Conjecture 4.1 in this case. In fact, this case also follows from the analogue of Lemma 2.7(b). ∎
@@
[…]
@@ 4.3. Verifying the spetsial conjecture
[…]